OpenStudy (♥ashley♥):

Is the number rational or irrational? Column A Column B 1. A. rational 3/7 B. irrational 2. --- 0.78 3. square root of 7 over 9 end root 4. 5 pi

OpenStudy (anonymous):

Pi is irr....

OpenStudy (anonymous):

1. a 2.b 3.b i think 4.? 5 b

OpenStudy (freckles):

if it can be written as an a/b where a and b are both integers (and a is not 0) then it is rational so this means any decimal that terminates or repeats is rational if it isn't rational and it is a real number, then it is irrational so things like pi, sqrt(2), sqrt(pi), sqrt(91), pi^2-1 are irrational numbers

OpenStudy (freckles):

why do you say .78 is irrational?

OpenStudy (anonymous):

cuz its a decimal

OpenStudy (freckles):

\[.78=\frac{78}{100}\]

OpenStudy (freckles):

decimals that terminate or repeat can be written as fractions when you have integer over integer

OpenStudy (freckles):

\[.12121212.... \text{ is also rational } \\ \text{ Let } x=.12121212.... \\100x=12.12121212.... \\ \text{ now subtract equations } \\ 100x-x=12.12121212.....-.1212121212.... \\ 99x=12 \\ x=\frac{12}{99}\] so .121212121212... can be written as 12/99 this again is an integer over an integer so therefore .12121212.... is rational